# Deciphering the Factors Influencing the Efficacy of Chain-of-Thought: Probability, Memorization, and Noisy Reasoning

### v1

@@ Abstract
Chain-of-Thought (CoT) prompting has been shown to enhance the multi-step reasoning capabilities of Large Language Models (LLMs). However, debates persist about whether LLMs exhibit abstract generalization or rely on shallow heuristics when given CoT prompts. To understand the factors influencing CoT reasoning we provide a detailed case study of the symbolic reasoning task of decoding shift ciphers, where letters are shifted forward some number of steps in the alphabet. GPT-4 achieves zero accuracy on most shift ciphers with standard prompting, but with CoT its accuracy improves to an average of 32%. By focusing on a single relatively simple task, we are able to identify three factors that systematically affect CoT performance: the probability of the task’s expected output (probability), what the model has implicitly learned during pre-training (memorization), and the number of intermediate operations involved in reasoning (noisy reasoning). We show that these factors can drastically influence the task accuracy; e.g., varying the output’s probability of occurrence can shift accuracy from $26\%\rightarrow 70\%$. We also demonstrate that it is essential for the model to explicitly produce intermediate steps as output that can be conditioned on to increase the probability of the correct answer. Our experiments indicate that as long as the model does so, the validity of the demonstrations in the prompt does not matter. Overall, we conclude that CoT prompting performance reflects both memorization and a probabilistic version of genuine reasoning. [^1] Preprint. Code and data are available at https://github.com/aksh555/deciphering_cot.
@@
Akshara Prabhakar1, Thomas L. Griffiths1,2, R. Thomas McCoy3,4 1Department of Computer Science, Princeton University 2Department of Psychology, Princeton University 3Department of Linguistics, Yale University 4Wu Tsai Institute, Yale University Correspondence: akshblr555@gmail.com
@@ 1 Introduction
Reasoning, one of the key aspects of human intelligence, is the process of thinking about something logically and systematically using evidence and past experiences to make a decision Wason (1968); Wason and Johnson-Laird (1972); Fagin et al. (2004). The impressive performance of Large Language Models (LLMs) across a wide range of tasks has spurred extensive research into their reasoning capabilities Huang and Chang (2022); Qiao et al. (2022). It remains unclear whether the behavior of these systems is based on true reasoning or other heuristics. Some results provide evidence that LLMs are able to reason Suzgun et al. (2022); Dasgupta et al. (2023); Saparov and He (2022), while others show that they still struggle on tasks that humans can easily solve through reasoning Han et al. (2022); Valmeekam et al. (2023); McCoy et al. (2023); Razeghi et al. (2022); Cao et al. (2023).
The Chain-of-Thought (CoT; Wei et al., 2022) prompting strategy has played a significant role in this debate. CoT involves prompting an LLM to generate a sequence of intermediate reasoning steps before producing the final answer, given a few in-context exemplar(s) of how to break the task into steps. CoT and its several variants Kojima et al. (2023); Zhou et al. (2022); Wang et al. (2023b) have been shown to substantially improve performance over standard prompting. Recent works have tried to identify which aspects of the demonstration contribute to CoT’s enhanced performance Huang and Chang (2022); Madaan and Yazdanbakhsh (2022); Jin et al. (2024), typically relying on assessing performance across a wide range of tasks.
@@
Figure 1: Overview. (1) Task: We have GPT-4 decode messages written in a shift cipher, in which each letter is shifted a fixed number of positions forward in the alphabet. (2) With standard prompting, GPT-4 performs poorly across most shift levels. (3) However, GPT-4 scores nearly perfectly on an isomorphic task based on numbers rather than letters. (4) With CoT prompting, GPT-4 adopts probabilistic and memorization-influenced noisy reasoning. That is, its performance (right) combines the trends we have hypothesized for each of the three factors on the left.
Our results suggest that CoT performance reflects three factors: probability, memorization, and noisy reasoning. First, the accuracy of CoT is affected by the probability of the correct output, with more probable outputs resulting in a stronger effect of CoT. Second, performance is higher when memorization is possible, as indicated by the frequency of encountering different shift cipher variants during pre-training. The effects of probability and memorization show that CoT performance is not fully systematic abstract reasoning. Nonetheless, CoT performance is not solely driven by superficial heuristics: it also shows some hallmarks of true reasoning—albeit a noisy version of true reasoning, in which the error rate increases along with task difficulty (where we quantify task difficulty by the number of implicit reasoning steps involved). In addition, we find evidence that the effect of CoT fundamentally depends on generating sequences of words that increase the probability of the correct answer when conditioned upon; as long as this is the case, CoT can thus succeed even when the demonstrations in the prompt are invalid. In the ongoing debate about whether LLMs reason or memorize, our results thus support a reasonable middle-ground: LLM behavior displays aspects of both memorization and reasoning, and also reflects the probabilistic origins of these models.
@@ 2 Related Work
It has been argued that LLMs can learn a task purely from demonstrations given in their context without any additional training Brown et al. (2020), a phenomenon known as in-context learning (ICL). There have been many investigations into how ICL operates. Theoretical frameworks have modeled the pretraining data as a mixture of Hidden Markov Models Xie et al. (2022) have argued that ICL is the result of implicit Bayesian inference Zhang et al. (2023), an argument supported with evidence from synthetic data and tasks Chiang and Yogatama (2024). The emergence of ICL has been attributed to factors including data distributional properties Chan et al. (2022), pretraining term frequencies Razeghi et al. (2022), and the creation of task vectors Hendel et al. (2023). However, the extent to which ICL is true learning is unclear Pan et al. (2023); Min et al. (2022). For example, Kossen et al. (2024) shows that ICL relies on in-context label information but cannot fully overcome preferences acquired during pre-training, which is evidence against the view that ICL is true learning.
@@
Theoretical arguments have been formulated about how CoT improves the expressivity Feng et al. (2023); Li et al. (2024) and sample complexity Li et al. (2023) of ICL. Empirical studies have shown that CoT performance can be dramatically influenced by many features of the CoT prompt Madaan and Yazdanbakhsh (2022); Wu et al. (2023); Jin et al. (2024); for example, the relevance and ordering of reasoning fragments is more important than their accuracy Wang et al. (2023a, b); Ye and Durrett (2022), and minor input perturbations can substantially bias models’ answers Turpin et al. (2024), indicating a lack of general reasoning ability Stechly et al. (2024).
@@ 3 Approach
Encoding a message with a shift cipher involves replacing every letter with another letter that is some fixed number of positions (called $shift\_level$) forward in the alphabet; decoding is the reverse (shifting backward) as shown in Figure 1. These are also known as rotation ciphers, since they rotate the alphabet forward some number of steps, and are given the name rot-$k$ where $k$ corresponds to $shift\_level$. For example, given the test word “FDW” and that rot-3 encryption has been used ($shift\_level=3$), decoding involves shifting every letter $3$ steps backward—i.e., F $\rightarrow$ C, D $\rightarrow$ A, and W $\rightarrow$ T to obtain “CAT” as the output. In our experiments, we give an LLM a single word encoded with a shift cipher and ask it to decode this text to recover the original word.
@@ 3.1 Motivation for using shift ciphers
Our main reason for using shift ciphers is because they involve a sharp dissociation between task complexity and task frequency (a key factor in memorization). The complexity of the decipherment task is determined by the shift level—ciphers that require more intermediate steps are more complex. Different shift levels also vary in their frequency in internet text, and hence in the training data of large language models. Specifically, rot-13 is widely used in internet forums to conceal text such as puzzle solutions and spoilers, and rot-3 and rot-1 commonly appear in tutorials on decipherment (rot-3 is also known as the Caesar cipher, having apparently been used by the eponymous Caesar to encrypt his messages). In addition, shift ciphers facilitate investigation of the effect of probability because the correct answer can be any string, allowing us to modulate the probability of that string easily. Further, the systematic nature of the task makes it easy to generate examples and to verify correctness. Finally, decoding each letter in the message is an independent step, allowing us to easily analyze these individual steps.
McCoy et al. (2023) previously evaluated LLMs on shift ciphers, focusing on standard prompting along with some initial results using CoT. We conduct a much more extensive investigation into LLM behavior when prompted with CoT.
@@ 3.2 The effect of CoT on shift ciphers
To provide an initial exploration of the effect of CoT and other aspects of prompting on shift ciphers, we constructed a dataset comprising 7-letter words having exactly 2 tokens (measured using the tokenizer used by GPT-4) to control for confounding factors relating to tokenization. The words were scored by their log probability and arranged in descending order. Subsequently, five bins were formed by selecting equidistant log probability values as centers, with bin1 having the highest probability and bin5 having the lowest probability. For each bin, we constructed a set of 150 words by choosing the words with log probability values nearest to each respective center (refer to Appendix A.1 for further details).
@@
  • Standard. This is a prompt with just the description of the task and demonstration but no reasoning steps. (Appendix A.2 Figure 5).
  • Text-based CoT (Text-CoT). This prompt encourages the model to decode a message one letter at a time (Figure 2). We chose this way of framing the CoT prompt following McCoy et al. (2023), which tried several variants and found this to be the best. To get a reasoning step correct, the model must have learned the alphabet during pre-training.
  • Math-based CoT (Math-CoT). The prompt (Appendix A.2 Figure 7) encourages a reasoning pipeline that involves translating each letter to a number, performing the shift by applying arithmetic to this number, then converting the result back to a letter. The prompt also specifies the mapping between letters and positions, eliminating the need for the model to have internalized the positions of the letters in the alphabet.
@@
We ran experiments using GPT-4 (gpt-4-0613 version) OpenAI (2024) because it is state-of-the-art in many areas. We set temperature to 00 and max_new_tokens to $200$. Using GPT-4 helps us to control several sources of extraneous errors making it easier to focus on the task itself and isolate the factors affecting CoT: It ensures that the format of the demonstration is closely followed, and that copy errors (errors in copying information from the prompt such as letters from encoded text and letter-position mappings) are rare.
Figure 1 provides some initial results. Using standard prompts, GPT-4 gets zero accuracy across most shift levels, but it improves substantially (to an average accuracy of 32%) when Text-CoT is used; this result replicates the finding in McCoy et al. (2023) that CoT is helpful for shift ciphers, but still remains far from perfect. However, with Number-CoT, GPT-4’s performance becomes nearly perfect (more details are in Appendix A.2).
@@ 4 Disentangling the factors influencing CoT performance
We consider four types of reasoning processes that GPT-4 might be adopting.
@@
    Noisy reasoning is like symbolic reasoning but with the addition of noise that introduces some possibility of each intermediate operation in a reasoning step being wrong. Thus, if the system uses noisy reasoning, we should see accuracy decrease as we increase the number of operations that need to be performed. Shift ciphers let us test this possibility: by varying $shift\_level$, we can modulate the number of operations that need to be performed in every reasoning step, and observe if accuracy varies accordingly.
@@
    Memorization is a strategy in which a system memorizes the tasks it has encountered in pre-training but does not generalize to new tasks. If memorization is all that GPT-4 does, we should see higher performance in the cases that are frequently encountered during pre-training than the ones that are not. McCoy et al. (2023) show that $13$ is by far the most common shift level in natural corpora because this shift level (sometimes called rot-13) is popular in some online communities. Thus, a hallmark of memorization would be much higher accuracy at $13$ than other shift levels.
@@
Figure 3 illustrates the hypothetical performance trends that would be observed in a system adopting each reasoning approach. These approaches are not mutually exclusive; e.g., a reasoner could be influenced by both probability and memorization. Indeed, as discussed in the next two paragraphs, we find that GPT-4’s performance when prompted with CoT displays hallmarks of several different types of reasoning (Figure 1, panel 4).
Figure 3: Hypothetical accuracy vs. shift-level for various types of reasoning. Under noisy one-way, the model only shifts letters backward; under noisy two-way, it adopts the shorter path between going forward and backward. The hypothetical memorization accuracy is based on shift level frequencies in internet corpora. Probabilistic would involve much higher scores on high probability than low probability.
First, accuracy generally decreases as the shift level increases, a hallmark of noisy reasoning. GPT-4’s performance is, in particular, indicative of a two-way version of noisy reasoning in which it can decode a message by shifting letters forward or backward (e.g., instead of decoding a shift of 25 by shifting back 25 letters, it could instead shift forward 1 letter, as doing so requires fewer steps); this two-way nature shows up in the way that accuracy increases as the shift level changes from 20 to 25.
Second, evidence of probabilistic reasoning can be seen in the fact that accuracy is substantially higher for bin 1 (the highest-probability bin) than bin 5 (the lowest-probability bin). Finally, although a shift level of 13 requires the most reasoning steps of any shift level (assuming decoding can be done forward or backward), there is a spike in accuracy at a shift level of 13. As discussed above, this spike is a hallmark of memorization since 13 is the most common shift level in natural corpora.

### v2

@@ Abstract
Chain-of-Thought (CoT) prompting has been shown to enhance the multi-step reasoning capabilities of Large Language Models (LLMs). However, debates persist about whether LLMs exhibit abstract generalization or rely on shallow heuristics when given CoT prompts. To understand the factors influencing CoT reasoning we provide a detailed case study of the symbolic reasoning task of decoding shift ciphers Andress (2014), where letters are shifted forward some number of steps in the alphabet. We analyze the pattern of results produced by three LLMs—GPT-4, Claude 3, and Llama 3.1—performing this task using CoT prompting. By focusing on a single relatively simple task, we are able to identify three factors that systematically affect CoT performance: the probability of the task’s expected output (probability), what the model has implicitly learned during pre-training (memorization), and the number of intermediate operations involved in reasoning (noisy reasoning). We show that these factors can drastically influence task accuracy across all three LLMs; e.g., when tested with GPT-4, varying the output’s probability of occurrence shifts accuracy from $26\%$ to $70\%$. Overall, we conclude that CoT prompting performance reflects both memorization and a probabilistic version of genuine reasoning. [^2] Code and data are available at https://github.com/aksh555/deciphering_cot.
@@
Akshara Prabhakar,1 Thomas L. Griffiths,1,2 R. Thomas McCoy3,4 thanks: Work begun while at Princeton University. 1Department of Computer Science, Princeton University 2Department of Psychology, Princeton University 3Department of Linguistics, Yale University 4Wu Tsai Institute, Yale University Correspondence: akshblr555@gmail.com
@@ 1 Introduction
Reasoning, one of the key aspects of human intelligence, is the process of thinking about something logically and systematically using evidence and past experiences to make a decision Wason (1968); Wason and Johnson-Laird (1972); Fagin et al. (2004). The impressive performance of Large Language Models (LLMs) across a wide range of tasks has spurred extensive research into their reasoning capabilities Huang and Chang (2023); Qiao et al. (2023). It remains unclear whether the behavior of these systems is based on true reasoning or on shallow heuristics. Some results provide evidence that LLMs are able to reason Suzgun et al. (2023); Lampinen et al. (2024); Saparov and He (2023), while others show that they still struggle on tasks that humans can easily solve via reasoning Han et al. (2022); Valmeekam et al. (2023); McCoy et al. (2023); Razeghi et al. (2022); Cao et al. (2024).
The Chain-of-Thought (CoT; Wei et al., 2022) prompting strategy has played a significant role in this debate. CoT involves prompting an LLM to generate a sequence of intermediate reasoning steps before producing the final answer, given some in-context exemplar(s) of how to break the task into steps. CoT and its several variants Kojima et al. (2022); Zhou et al. (2023); Wang et al. (2023b) have been shown to substantially improve performance over standard prompting. Recent works have tried to identify which aspects of the demonstration contribute to CoT’s enhanced performance Huang and Chang (2023); Madaan and Yazdanbakhsh (2022); Jin et al. (2024), typically relying on assessing performance across a wide range of tasks.
@@
Figure 1: Overview. (1) Task: We have LLMs decode messages written in a shift cipher, in which each letter is shifted a fixed number of positions forward in the alphabet. (2) With standard prompting, GPT-4 performs poorly across most shift levels. (3) However, GPT-4 scores nearly perfectly on an isomorphic task based on numbers rather than letters. (4) With CoT prompting, GPT-4 adopts probabilistic and memorization-influenced noisy reasoning. That is, its performance (right) combines the trends we have hypothesized for each of the three factors on the left.
Our results suggest that CoT performance reflects three factors: probability, memorization, and noisy reasoning. First, the accuracy of CoT is affected by the probability of the correct output, with more probable outputs resulting in a stronger effect of CoT. Second, performance is higher when memorization is possible, as indicated by the frequency of encountering different shift cipher variants during pre-training. The effects of probability and memorization show that CoT performance is not fully systematic abstract reasoning. Nonetheless, CoT performance is not solely driven by superficial heuristics: it also shows some hallmarks of true reasoning—albeit a noisy version of true reasoning, in which the error rate increases along with task difficulty (where we quantify task difficulty by the number of implicit reasoning steps involved). In addition, we find evidence that the effect of CoT fundamentally depends on generating sequences of words that increase the probability of the correct answer when conditioned upon; as long as this is the case, CoT can thus succeed even when the demonstrations in the prompt are invalid. In the ongoing debate about whether LLMs reason or memorize Feldman (2020); Zhang et al. (2023a); Magar and Schwartz (2022); Srivastava et al. (2024); Antoniades et al. (2024), our results thus support a reasonable middle-ground: LLM behavior displays aspects of both memorization and reasoning, and also reflects the probabilistic origins of these models.
@@ 2 Related Work
It has been argued that LLMs can learn a task purely from demonstrations given in their context without any additional training Brown et al. (2020), a phenomenon known as in-context learning (ICL). There have been many investigations into how ICL operates. Theoretical frameworks have modeled the pretraining data as a mixture of Hidden Markov Models Xie et al. (2022) and have argued that ICL is the result of implicit Bayesian inference Zhang et al. (2023b), an argument supported with evidence from synthetic data and tasks Chiang and Yogatama (2024). The emergence of ICL has been attributed to factors including data distributional properties Chan et al. (2022), pretraining term frequencies Razeghi et al. (2022), and the creation of task vectors Hendel et al. (2023). However, the extent to which ICL is true learning is unclear Pan et al. (2023); Min et al. (2022). For example, Kossen et al. (2024) shows that ICL relies on in-context label information but cannot fully overcome preferences acquired during pre-training, which is evidence against the view that ICL is true learning.
@@
Theoretical arguments have been formulated about how CoT improves the expressivity Feng et al. (2023); Li et al. (2024); Merrill and Sabharwal (2024) and sample complexity Li et al. (2023) of ICL. Empirical studies have shown that CoT performance can be dramatically influenced by many features of the CoT prompt Madaan and Yazdanbakhsh (2022); Wu et al. (2023); Jin et al. (2024); for example, the relevance and ordering of reasoning fragments is more important than their accuracy Wang et al. (2023a, b); Ye and Durrett (2022), and minor input perturbations can substantially bias models’ answers Turpin et al. (2024), suggesting that the models lack general reasoning abilities Stechly et al. (2024).
@@ 3 Approach
Encoding a message with a shift cipher involves replacing every letter with another letter that is some fixed number of positions (called $shift\_level$) forward in the alphabet; decoding is the reverse (shifting backward) as shown in Figure 1. These are also known as rotation ciphers since they rotate the alphabet forward some number of steps, and they are given the name rot-$k$ where $k$ corresponds to $shift\_level$. For example, given the test word “FDW” and that rot-3 encryption has been used ($shift\_level=3$), decoding involves shifting every letter $3$ steps backward—i.e., F $\rightarrow$ C, D $\rightarrow$ A, and W $\rightarrow$ T to obtain “CAT” as the output. In our experiments, we give an LLM a single word encoded with a shift cipher and ask it to decode this text to recover the original word.
@@ 3.1 Motivation for using shift ciphers
Our main reason for using shift ciphers is because they involve a sharp dissociation between task complexity and task frequency (a key factor in memorization). The complexity of the decipherment task is determined by the shift level—ciphers that require more intermediate steps are more complex. Different shift levels also vary in their frequency in internet text McCoy et al. (2023), and hence in the training data of large language models. Specifically, rot-13 is widely used in internet forums to conceal text such as puzzle solutions and spoilers, and rot-3 and rot-1 commonly appear in tutorials on decipherment (rot-3 is also known as the Caesar cipher, having apparently been used by the eponymous Caesar to encrypt his messages). In addition, shift ciphers facilitate investigation of the effect of probability because the correct answer can be any string, allowing us to modulate the probability of that string easily. Further, the systematic nature of the task makes it easy to generate examples and to verify correctness. Finally, decoding each letter in the message is an independent step, allowing us to easily analyze these individual steps.
McCoy et al. (2023) previously evaluated GPT models on shift ciphers, focusing on standard prompting along with some initial results using CoT. They study the effect of only probability and memorization, while we conduct a more extensive investigation into LLM behavior when prompted with CoT by additionally studying the influence of complexity and analyzing more models. Importantly, we add nuance to their findings by arguing for a middle-ground viewpoint that acknowledges the LLM weaknesses identified by McCoy et al. (2023) but also brings in novel observations that highlight the hallmarks of true reasoning that are present in these systems.
@@ 3.2 The effect of CoT on shift ciphers
We constructed a dataset comprising 7-letter words having exactly 2 tokens (measured using the tokenizer used by GPT-4) to control for confounding factors relating to tokenization. We found all 3-letter and 4-letter tokens from the lowercase English alphabet and formed words by considering possible combinations of 3-letter word-initial tokens followed by 4-letter non-word-initial tokens. Following McCoy et al. (2023), we compute the log probability as the log probability that GPT-2 Radford et al. (2019) assigns to the sentence ‘The word is "WORD"’, minus the log probability that it assigns to ‘The word is "’; thus, this yields the log probability assigned to just the word and the following quotation mark in the context of ‘The word is "’. The closing quotation mark is included because it indicates the end of the word. The words were scored by their log probability and arranged in descending order. Subsequently, five bins were formed by selecting equidistant log probability values as centers, with bin1 having the highest probability and bin5 having the lowest probability. We manually checked the words in this dataset and filtered them to ensure there were no inappropriate words used to obtain 150 words for each bin. We partitioned the 150 examples into two subsets: a subset containing 100 words used to evaluate GPT-4, and a subset containing 50 words used to evaluate logistic regression models that were fitted to GPT-4’s performance on the 100-word subset. We prepared the inputs for the models by producing the shift-cipher-encoded versions of the words from the 5 probability bins across 25 shift levels (1 to 25). We ran all evaluations a single time; the accuracies that we report are accuracies over these 100-example sets.
@@
  • Standard. This is a prompt with just the description of the task and demonstration but no reasoning steps (Figure 2).
    Rot-13 text: "fgnl"
    Original text: "stay"
  • Text-CoT. This prompt encourages the model to decode a message one letter at a time (Figure 3). We chose this way of framing the CoT prompt following McCoy et al. (2023), who tried several variants and found this to be the best. To get a reasoning step correct, the model must have learned the alphabet during pre-training.
  • Math-CoT. The prompt (Appendix A.1 Figure 9) encourages a reasoning pipeline that involves translating each letter to a number, performing the shift by applying arithmetic to this number, then converting the result back to a letter. The prompt also specifies the mapping between letters and positions, eliminating the need for the model to have internalized the positions of the letters in the alphabet.
@@
We ran experiments using both open and closed source models: GPT-4 (gpt-4-0613) OpenAI (2023), Claude 3 (claude-3-opus-20240229) Anthropic (2024), and Llama-3.1-405B-Instruct MetaAI (2024). The reason for using such strong models is that their shift cipher performance is significantly improved by prompting with chain of thought. Additionally, this helps us to control several sources of extraneous errors making it easier to focus on the task itself and isolate the factors affecting CoT: It ensures that the format of the demonstration is closely followed, and that copy errors (errors in copying information from the prompt such as letters from encoded text and letter-position mappings) are rare. We set temperature to 00 and max_new_tokens to $200$.
Figure 1 provides some initial results for GPT-4. Using standard prompts, GPT-4 gets zero accuracy across most shift levels, but it improves substantially (to an average accuracy of 32%) when Text-CoT is used; this result replicates the finding in McCoy et al. (2023) that CoT is helpful for shift ciphers but still remains far from perfect. However, with Number-CoT, GPT-4’s performance becomes nearly perfect (more details are in Appendix A.1).
@@ 4 Disentangling the factors influencing CoT performance
We consider four types of reasoning processes that LLMs might be adopting.
@@
    Noisy reasoning is like symbolic reasoning but with the addition of noise that introduces some possibility of each intermediate operation in a reasoning step being wrong. Thus, if the system uses noisy reasoning, we should see accuracy decrease as we increase the number of operations that need to be performed. Shift ciphers let us test this possibility: by varying $shift\_level$, we can modulate the number of operations that need to be performed in every reasoning step and observe if accuracy varies accordingly.
@@
    Memorization is a strategy in which a system memorizes the tasks it has encountered in pre-training but does not generalize to new tasks. If memorization is all that LLMs do, we should see higher performance in the cases that are frequently encountered during pre-training than the ones that are not. McCoy et al. (2023) show that $13$ is by far the most common shift level in natural corpora because this shift level (sometimes called rot-13) is popular in some online communities. Thus, a hallmark of memorization would be much higher accuracy at $13$ than other shift levels.
@@
Figure 4 illustrates the hypothetical performance trends that would be observed in a system adopting each reasoning approach. These approaches are not mutually exclusive; e.g., a reasoner could be influenced by both probability and memorization. Indeed, as discussed below, we find that LLMs’ performance when prompted with CoT displays hallmarks of several different types of reasoning; see Figure 1, panel 4 for GPT-4 and Figure 11 in the Appendix for Claude 3 and Llama 3.1.
Figure 4: Hypothetical accuracy vs. shift-level for various types of reasoning. Under noisy one-way, the model only shifts letters backward; under noisy two-way, it adopts the shorter path between going forward and backward. The hypothetical memorization accuracy is based on shift level frequencies in internet corpora. Probabilistic would involve much higher scores on high prob than low prob.
First, accuracy generally decreases as the shift level increases, a hallmark of noisy reasoning. LLMs’ performance is, in particular, indicative of a two-way version of noisy reasoning in which it can decode a message by shifting letters forward or backward (e.g., instead of decoding a shift of 25 by shifting back 25 letters, it could instead shift forward 1 letter, as doing so requires fewer steps); this two-way nature shows up in the way that accuracy increases as the shift level changes from 20 to 25.
Second, evidence of probabilistic reasoning can be seen in the fact that accuracy is substantially higher for high prob (the highest-probability bin, bin 1) than low prob (the lowest-probability bin, bin 5). High prob / low prob refer to the probability of the words that are the correct answers when our examples are decoded, where probability is quantified using GPT-2 as described in Section 3.2. The “high prob” cases are common words such as {’mariner’, ’shrines’, ’paywall’, …}, while the “low prob” cases are nonsense letter sequences such as {’xcbrouw’, ’jsxrouw’, ’levjspx’, …}.
Finally, although a shift level of 13 requires the most reasoning steps of any shift level (assuming decoding can be done forward or backward), there is a spike in accuracy at a shift level of 13. As discussed above, this spike is a hallmark of memorization since 13 is the most common shift level in natural corpora.
For the upcoming detailed analysis experiments, we use GPT-4 as the main reference model, since Claude 3 and Llama 3.1 exhibited similar trends as GPT-4 on the evaluations presented so far.***The detailed results of Llama-3.1-405B-Instruct and claude-3-opus-20240229 are shown in Appendix §A.2.